# Controllability of the Fisher-Stefan system

### v1

@@ 2.1 Linearization
$\mathbf{P}:=\left\{(w,\theta)\in C^{2}([-1,1]\times[0,T])\times C^{1}([0,T])\;:\;w(-1,\cdot)=0\;\mbox{and}\;\right.$
$\left.\quad\quad\quad\quad w(1,\cdot)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,\cdot)w(x,\cdot)\mathrm{d}x+2\mu\theta\quad\mbox{in}\;(0,T)\right\}.$
@@
$\mathbf{B}((w,\theta),(\hat{w},\hat{\theta}))$ $:=\iint_{\tilde{Q}}\rho_{0}^{-2}\textbf{L}_{1}^{\star}(w,\theta)\textbf{L}^{\star}_{1}(\hat{w},\hat{\theta})\mathrm{d}x\mathrm{d}t+\int_{0}^{T}\rho_{0}^{-2}\textbf{L}_{2}^{\star}(w,\theta)\textbf{L}^{\star}_{2}(\hat{w},\hat{\theta})\mathrm{d}t$
@@
where $\textbf{L}_{1}^{\star}$ and $\textbf{L}_{2}^{\star}$ are defined in (2.2). We also define the linear form $\mathbf{F}:\mathbf{P}\longrightarrow\mathbb{R}$ by
@@
We claim that, $\mathbf{B}$ is a scalar product in $\mathbf{P}$ and $\mathbf{F}$ is continuous for the norm $\|\cdot\|_{\mathbf{B}}$ associated with the scalar product $\mathbf{B}$. Indeed, due to Carleman’s estimate (2.10) and Remark 3, there exists a constant $C:=C(s,\lambda,T)>0$ such that for all $(w,\theta)\in\mathbf{P}$, one has
$\iint_{\tilde{Q}}\rho^{-2}|w|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}\rho^{-2}|\theta|^{2}\mathrm{d}t+\int_{-1}^{1}|w(x,0)|^{2}\mathrm{d}x+|\theta(0)|^{2}\leq C\,\mathbf{P}((w,\theta),(w,\theta)).$ (2.17)
In particular, $\mathbf{B}$ is a scalar product in $\mathbf{P}$. To ensure the continuity of $\mathbf{F}$, using the Cauchy-Schwarz inequality and (2.17), we obtain
$|\mathbf{F}(w,\theta)|\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}+\|\rho g_{0}\|_{L^{2}(0,T)}+\|z_{0}\|_{L^{2}(-1,1)}+|k_{0}|\right)\|(w,\theta)\|_{\mathbf{B}}.$ (2.18)
@@
$\mathbf{B}(w^{*},\theta^{*}),(w,\theta))=\mathbf{F}(w,\theta)\quad\forall(w,\theta)\in\overline{\mathbf{P}}.$ (2.19)
Using (2.18) and (2.19), we obtain
$\|(w^{*},\theta^{*})\|_{\mathbf{B}}\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}+\|\rho g_{0}\|_{L^{2}(0,T)}+\|z_{0}\|_{L^{2}(-1,1)}+|k_{0}|\right).$ (2.20)
@@
$(z^{*},k^{*}):=\rho_{0}^{-2}(\mathbf{L}_{1}^{\star}(w^{*},\theta^{*}),\mathbf{L}_{2}^{\star}(w^{*},\theta^{*})),\quad v^{*}:=-\rho_{1}^{-2}w^{*}\mathds{1}_{\omega}.$ (2.21)
According to (2.21) and the definition of $\mathbf{B}$, we obtain
@@
Using this equation and estimate (2.20), we deduce estimate (2.15). As a consequence $(z^{*},k^{*})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, $v^{*}\in L^{2}(\omega\times(0,T))$ and from (2.19), $(z^{*},k^{*})$ is the unique solution by transposition of (2.1) with the control $v^{*}$. Moreover, due to the uniqueness of solutions by transposition of (2.1) and the regularity of initial data and term sources, we obtain $(z^{*},k^{*})$ is the unique strong solution of (2.1).
Now, it is a matter of proving estimate (2.16), to do this, let us set $(z,k):=\rho_{2}(z^{*},k^{*})$. Then $(z,k)$ is the solution of system:
$&\mathbf{L}_{1}(z,k)=\overline{h}\rho_{2,t}z^{*}+\rho_{2}f_{0}+\mathds{1}_{\omega}\rho_{2}v^{*}:=f^{*}&&\text{in}\;\tilde{Q},\\ &z(-1,\cdot)=z(1,\cdot)=0&&\text{in}\;(0,T),\\ &\mathbf{L}_{2}(z,k)=\rho_{2,t}k^{*}+\rho_{2}g_{0}:=g^{*}&&\text{in}\;(0,T),\\ &k(0)=\rho_{2}(0)k_{0},\\ &z(\cdot,0)=\rho_{2}(0)z_{0}&&\text{in}\;(-1,1),\end{aligned}\right.$ (2.22)
@@
$|\rho_{2,t}|\leq C\rho_{0},\quad\rho_{2}\leq C\rho_{1},\quad\rho_{1}\leq C\rho_{0}$ (2.23)
and estimate (2.15), we obtain $(f^{*},g^{*})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. Given that $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$, then $(z,k)\in H_{0}^{2,1}(\tilde{Q})\times H^{1}(0,T)$ is the strong solution of (2.22). Finally, from (2.3), (2.23) and (2.15), we obtain (2.16). ∎
@@
The proof of Proposition 2.5 allows for identifying a null control of system (2.1) by solving a variational problem, providing a numerical approach based on finite elements [27, 28, 29, 30]. However, approximating the space $P$ is highly challenging and computationally expensive. To overcome this difficulty, we chose to solve the controllability problem using PINNs (see Section 4), which enables direct handling of the nonlinear problem, unlike classical methods such as finite differences or finite elements.
3 Study of the nonlinear system (1.11)
We will establish the local null controllability of system (1.11) using Lyusternik-Graves’ Theorem:
Theorem 3.1 (Lyusternik-Graves’ Theorem).
Let $\mathds{X}$ and $\mathds{Y}$ be Banach spaces, and let $\Lambda:B_{\mathds{X}}(0,r)\subset\mathds{X}\rightarrow\mathds{Y}$ be a continuously differentiable mapping ($B_{\mathds{X}}(0,r)$ denote the ball in the space $\mathds{X}$ centered at 00 with radius $r$). Let us assume that the derivative $\Lambda^{\prime}(0):\mathds{X}\rightarrow\mathds{Y}$ is onto and let us set $\xi_{0}=\Lambda(0)$. Then there exist $\varepsilon>0$, a mapping $W:B_{\mathds{Y}}(\xi_{0},\varepsilon)\subset\mathds{Y}\rightarrow\mathds{X}$ and a constant $C>0$ satisfying:
@@
We then introduce the following appropriate spaces and mapping so that the conditions of Theorem 3.1 are satisfied:
@@
where $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ are defined in (2.2). These spaces are naturally equipped with the following norms:
@@
$+\left.\iint_{\tilde{Q}}\rho^{2}(t)|z_{x}(1,t)|^{2}|\tilde{z}_{x}(x,t)|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}\rho^{2}(t)|z(x,t)|^{2}|\tilde{z}(x,t)|^{2}\mathrm{d}x\mathrm{d}t\right).$ (3.1)
We apply the fact that $\rho_{2}k\in H^{1}(0,T)$ and the continuous embedding $H^{1}(0,T)\hookrightarrow L^{\infty}(0,T)$ for the first integral on the right-hand side of (3.1), the Sobolev continuous embedding $H^{2,1}(\tilde{Q})\hookrightarrow C([0,T];H^{1}(-1,1))$, the continuity of the normal derivative: $|z_{x}(1,t)|\leq C\|z(\cdot,t)\|_{H^{2}(-1,1)}$ for the second one and the fact that $H^{1}(-1,1)$ is a Banach algebra for third one, we obtain
@@
Due to this lemma and null controllability result of inhomogeneous linearized system (2.1) obtained in Proposition 2.5, $\Lambda^{\prime}(0,0,0)$ is surjective. We can therefore apply Lyusternik-Graves’ Theorem 3.1 to this data, in particular, for any sufficiently small $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$, there exists $(z,k,v)\in\mathds{X}$ such that $\Lambda(z,k,v)=(0,0,z_{0},k_{0})$. Consequently, $(z,k)$ is the solution of (2.4) associated to the control $v$, and the exponential growth of the weight $\rho_{0}$ as $t\rightarrow T^{-}$ and $\rho_{0}(z,k)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, ensures that
@@
This leads to the proof of Theorem 1.3 and, as explained in the introduction, we immediately obtain the main Theorem 1.2 (Note that the regularity of the boundary control is justified by the trace theorem [31, Theorem 2.1]).
@@ 4 Numerical Modeling with PINNs
In recent years, Physics-Informed-Neural-Networks (PINNs) have attracted considerable attention for solving mathematical models based on nonlinear PDEs; see, for example, [32, 33] and the references therein. This section addresses the numerical resolution of problem (1.5), and its linearized counterpart, by using Physics-Informed-Neural-Networks (PINNs).
@@ 4.1 Description of the numerical algorithm
For the sake of clarity in the exposition, let us consider the linearized system (2.1), with $\overline{\psi}=0,\overline{h}=1$ (which corresponds to Experiment 1 below), the nonlinear case being completely analogous. More precisely, we consider the system
{$&z_{t}-z_{xx}-z=\mathds{1}_{\omega}v&&\text{in}\;\Omega_{T}:=\Omega\times(0,T),\\ &z=0&&\text{on}\;\Gamma_{T}:=\partial\Omega\times(0,T),\\ &k^{\prime}+2\mu z_{x}(1,\cdot)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;\Omega,\end{aligned}\right.$
@@
$\begin{array}[]{lll}\mathcal{L}_{\text{int},PDE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int, }\Omega_{T}}\right)&=\displaystyle\sum_{j=1}^{N_{\text{int, }\Omega_{T}}}w_{j,\text{int, }\Omega_{T}}|\hat{z}_{t}(\boldsymbol{x}_{j})-\hat{z}_{xx}(\boldsymbol{x}_{j})-\hat{z}(\boldsymbol{x}_{j})-\mathds{1}_{\omega}\hat{v}(\boldsymbol{x}_{j})|^{2},&\boldsymbol{x}_{j}\in\mathcal{T}_{\text{int, }\Omega_{T}},\\ \\ \mathcal{L}_{\text{int },ODE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int,}T}\right)&=\displaystyle\sum_{j=1}^{N_{\text{int, }T}}w_{j,\text{int, }T}|\hat{k}^{\prime}(\boldsymbol{x}_{j})+2\mu\hat{z}_{x}(1,\boldsymbol{\tilde{x}}_{j})|^{2},&(\boldsymbol{x}_{j},\boldsymbol{\tilde{x}}_{j})\in\mathcal{T}_{\text{int, }T}\times\mathcal{T}_{\text{int, }\Omega_{T}},\\ \\ \mathcal{L}_{\Gamma}\left(\boldsymbol{\theta};\mathcal{T}_{\Gamma}\right)&=\displaystyle\sum_{j=1}^{N_{b}}w_{j,b}|\hat{z}(\boldsymbol{x}_{j};\boldsymbol{\theta})|^{2},&\boldsymbol{x}_{j}\in\mathcal{T}_{\Gamma},\\ \\ \mathcal{L}_{t=0}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=0}\right)&=\displaystyle\sum_{j=1}^{N_{0}}w_{j,0}|\hat{z}(\boldsymbol{x}_{j};\boldsymbol{\theta})-z_{0}(\boldsymbol{x}_{j})|^{2},&\boldsymbol{x}_{j}\in\mathcal{T}_{t=0},\\ \\ \mathcal{L}_{t=0}^{k}\left(\boldsymbol{\theta}\right)&=|\hat{k}(0;\boldsymbol{\theta})-k_{0}|,\\ \\ \mathcal{L}_{t=T}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=T}\right)&=\displaystyle\sum_{j=1}^{N_{T}}w_{j,T}|\hat{z}(\boldsymbol{x}_{j};\boldsymbol{\theta})|^{2},&\boldsymbol{x}_{j}\in\mathcal{T}_{t=T},\\ \\ \mathcal{L}_{t=T}^{k}\left(\boldsymbol{\theta}\right)&=|\hat{k}(T;\boldsymbol{\theta})|,&\\ \\ \mathcal{L}_{\hat{z}\geq 0}(\boldsymbol{\theta})&=\displaystyle\sum_{j=1}^{N_{\text{int, }\Omega_{T}}}w_{j,\text{int, }\Omega_{T}}\frac{1}{\beta}\log\left(1+e^{-\beta\hat{z}(\boldsymbol{x}_{j})}\right)|^{2},&\boldsymbol{x}_{j}\in\mathcal{T}_{\text{int, }\Omega_{T}}.\end{array}$
@@
$\begin{array}[]{ll}\mathcal{L}\left(\boldsymbol{\theta};\mathcal{T}\right)&=w_{1}\mathcal{L}_{\text{int},PDE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int, }\Omega_{T}}\right)+w_{2}\mathcal{L}_{\text{int },ODE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int, }T}\right)\\ &+w_{3}\mathcal{L}_{\Gamma}\left(\boldsymbol{\theta};\mathcal{T}_{\Gamma}\right)\\ &+w_{4}\mathcal{L}_{t=0}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=0}\right)+w_{5}\mathcal{L}_{t=0}^{k}\left(\boldsymbol{\theta}\right)\\ &+w_{6}\mathcal{L}_{t=T}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=T}\right)+w_{7}\mathcal{L}_{t=T}^{k}\left(\boldsymbol{\theta}\right)\\ &+w_{8}\mathcal{L}_{\hat{z}\geq 0}(\boldsymbol{\theta}),\end{array}$ (4.1)
@@
$\hat{z}(x,t)$$\hat{k}(t)$$\hat{v}(x,t)$$\hat{z}_{t}-\hat{z}_{xx}-\hat{z}-\mathds{1}_{\omega}\hat{v}=0$$\hat{k}^{\prime}+2\mu\hat{z}_{x}(1,\cdot)=0$$\hat{z}(-1,t)=\hat{z}(1,t)=0$$\hat{z}(x,0)=\hat{z}_{0}(x)$$\hat{k}(0)=\hat{k}_{0}$$\hat{z}(x,T)=\hat{z}_{T}(x)$$\hat{k}(T)=\hat{k}_{T}$$\min_{\theta}\mathcal{L}(\hat{z},\hat{k},\hat{v};\theta)$$\theta^{\ast}$$\hat{z}(x,t;\theta^{\ast})$$\hat{k}(t,\theta^{\ast})$$\hat{v}(x,t;\theta^{\ast})$
@@
Step 4: Training process. The final step in the PINN algorithm amounts to minimizing (4.1), i.e.,

### v2

@@ 2.3 Null controllability of system (14)
$\mathbf{P}:=\left\{(w,\theta)\in C^{2}([-1,1]\times[0,T])\times C^{1}([0,T])\;:\;w(-1,t)=0\;\mbox{and}\;\right$.
$\left.\quad\quad\quad\quad w(1,t)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,t)w(x,t)\mathrm{d}x+2\mu\theta(t)\quad\mbox{in}\;(0,T)\right\}.$
@@
$)\displaystyle\mathbf{B}((w,\theta),(\hat{w},\hat{\theta}$)) $:=\iint_{\tilde{Q}}\rho_{0}^{-2}\textbf{L}_{1}^{\star}(w,\theta)\textbf{L}^{\star}_{1}(\hat{w},\hat{\theta})\mathrm{d}x\mathrm{d}t+\int_{0}^{T}\rho_{0}^{-2}\textbf{L}_{2}^{\star}(w,\theta)\textbf{L}^{\star}_{2}(\hat{w},\hat{\theta})\mathrm{d}t$
@@
where $\textbf{L}_{1}^{\star}$ and $\textbf{L}_{2}^{\star}$ are defined in (15). We also define the linear form $\mathbf{F}:\mathbf{P}\longrightarrow\mathbb{R}$ by
@@
We claim that, $\mathbf{B}$ is a scalar product in $\mathbf{P}$ and $\mathbf{F}$ is continuous for the norm $\|\cdot\|_{\mathbf{B}}$ associated with the scalar product $\mathbf{B}$. Indeed, due to Carleman’s estimate (23) and Remark 4, there exists a constant $C:=C(s,\lambda,T)>0$ such that for all $(w,\theta)\in\mathbf{P}$, one has
$\iint_{\tilde{Q}}\rho^{-2}|w|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}\rho^{-2}|\theta|^{2}\mathrm{d}t+\int_{-1}^{1}|w(x,0)|^{2}\mathrm{d}x+|\theta(0)|^{2}\leq C\,\mathbf{P}((w,\theta),(w,\theta)).$ (30)
In particular, $\mathbf{B}$ is a scalar product in $\mathbf{P}$. To ensure the continuity of $\mathbf{F}$, using the Cauchy-Schwarz inequality and (30), we obtain
$|\mathbf{F}(w,\theta)|\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}+\|\rho g_{0}\|_{L^{2}(0,T)}+\|z_{0}\|_{L^{2}(-1,1)}+|k_{0}|\right)\|(w,\theta)\|_{\mathbf{B}}.$ (31)
@@
$\mathbf{B}(w^{*},\theta^{*}),(w,\theta))=\mathbf{F}(w,\theta)\quad\forall(w,\theta)\in\overline{\mathbf{P}}.$ (32)
Using (31) and (32), we obtain
$\|(w^{*},\theta^{*})\|_{\mathbf{B}}\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}+\|\rho g_{0}\|_{L^{2}(0,T)}+\|z_{0}\|_{L^{2}(-1,1)}+|k_{0}|\right).$ (33)
@@
$(z^{*},k^{*}):=\rho_{0}^{-2}(\mathbf{L}_{1}^{\star}(w^{*},\theta^{*}),\mathbf{L}_{2}^{\star}(w^{*},\theta^{*})),\quad v^{*}:=-\rho_{1}^{-2}w^{*}\mathds{1}_{\omega}.$ (34)
According to (34) and the definition of $\mathbf{B}$, we obtain
@@
Using this equation and estimate (33), we deduce estimate (28). As a consequence $(z^{*},k^{*})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, $v^{*}\in L^{2}(\omega\times(0,T))$ and from (32), $(z^{*},k^{*})$ is the unique solution by transposition of (14) with the control $v^{*}$. Moreover, due to the uniqueness of solutions by transposition of (14) and the regularity of initial data and term sources, we obtain $(z^{*},k^{*})$ is the unique strong solution of (14).
Now, it is a matter of proving estimate (29), to do this, let us set $(z,k):=\rho_{2}(z^{*},k^{*})$. Then $(z,k)$ is the solution of system:
$\left\{\begin{aligned} &\mathbf{L}_{1}(z,k)=\overline{h}\rho_{2,t}z^{*}+\rho_{2}f_{0}+\mathds{1}_{\omega}\rho_{2}v^{*}:=f^{*}&&\text{in}\;\tilde{Q},\\ &z(-1,t)=z(1,t)=0&&\text{in}\;(0,T),\\ &\mathbf{L}_{2}(z,k)=\rho_{2,t}k^{*}+\rho_{2}g_{0}:=g^{*}&&\text{in}\;(0,T),\\ &k(0)=\rho_{2}(0)k_{0},\\ &z(\cdot,0)=\rho_{2}(0)z_{0}&&\text{in}\;(-1,1),\end{aligned}\right$. (35)
@@
$|\rho_{2,t}|\leq C\rho_{0},\quad\rho_{2}\leq C\rho_{1},\quad\rho_{1}\leq C\rho_{0}$ (36)
and estimate (28), we obtain $(f^{*},g^{*})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. Given that $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$, then $(z,k)\in H_{0}^{2,1}(\tilde{Q})\times H^{1}(0,T)$ is the strong solution of (35). Finally, from (16), (36) and (28), we obtain (29). ∎
@@
The proof of Proposition 6 allows for identifying a null control of system (14) by solving a variational problem, providing a numerical approach based on finite elements [15, 17, 19, 33]. However, approximating the space $P$ is highly challenging and computationally expensive. To overcome this difficulty, we chose to solve the controllability problem using PINNs (see Section 4), which enables direct handling of the nonlinear problem, unlike classical methods such as finite differences or finite elements.
3 Study of the nonlinear system (11)
We will establish the local null controllability of system (11) using Lyusternik-Graves’ Theorem:
Theorem 7 (Lyusternik-Graves’ Theorem).
Let $\mathds{X}$ and $\mathds{Y}$ be Banach spaces, and let $\Lambda:B_{\mathds{X}}(0,r)\subset\mathds{X}\rightarrow\mathds{Y}$ be a continuously differentiable mapping ($B_{\mathds{X}}(0,r)$ denote the ball in the space $\mathds{X}$ centered at 0 with radius $r$). Let us assume that the derivative $\Lambda^{\prime}(0):\mathds{X}\rightarrow\mathds{Y}$ is onto and let us set $\xi_{0}=\Lambda(0)$. Then there exist $\varepsilon>0$, a mapping $W:B_{\mathds{Y}}(\xi_{0},\varepsilon)\subset\mathds{Y}\rightarrow\mathds{X}$ and a constant $C>0$ satisfying:
@@ 3 Study of the nonlinear system (11)
We then introduce the following appropriate spaces and mapping so that the conditions of Theorem 7 are satisfied:
@@
where $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ are defined in (15). These spaces are naturally equipped with the following norms:
@@
$+\left.\iint_{\tilde{Q}}\rho^{2}(t)|z_{x}(1,t)|^{2}|\tilde{z}_{x}(x,t)|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}\rho^{2}(t)|z(x,t)|^{2}|\tilde{z}(x,t)|^{2}\mathrm{d}x\mathrm{d}t\right).$ (37)
We apply the fact that $\rho_{2}k\in H^{1}(0,T)$ and the continuous embedding $H^{1}(0,T)\hookrightarrow L^{\infty}(0,T)$ for the first integral on the right-hand side of (37), the Sobolev continuous embedding $H^{2,1}(\tilde{Q})\hookrightarrow C([0,T];H^{1}(-1,1))$, the continuity of the normal derivative: $|z_{x}(1,t)|\leq C\|z(\cdot,t)\|_{H^{2}(-1,1)}$ for the second one and the fact that $H^{1}(-1,1)$ is a Banach algebra for third one, we obtain
@@
Due to this lemma and null controllability result of inhomogeneous linearized system (14) obtained in Proposition 6, $\Lambda^{\prime}(0,0,0)$ is surjective. We can therefore apply Lyusternik-Graves’ Theorem 7 to this data, in particular, for any initial data $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$ with sufficiently small norm
$\|(z_{0},k_{0})\|_{H^{1}_{0}(-1,1)\times\mathbb{R}}<\varepsilon,$
for some sufficiently small $\varepsilon>0$, there exists $(z,k,v)\in\mathds{X}$ such that
$\Lambda(z,k,v)=(0,0,z_{0},k_{0}).$
Consequently, $(z,k)$ is the solution of (17) associated to the control $v$, and the exponential growth of the weight $\rho_{0}$ as $t\rightarrow T^{-}$ and $\rho_{0}(z,k)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, ensures that
@@
Moreover, we have
$\|(z,k,v)\|_{\mathds{X}}\leq C(\|(z_{0},k_{0})\|_{H^{1}(-1,1)\times\mathbb{R}},$
for some constant $C>0$ independent of $\varepsilon$. This leads to the proof of Theorem 2 and, as explained in the introduction and Remark 2, we immediately obtain the main Theorem 1.
@@ 4 Numerical Modeling with PINNs
In recent years, Physics-Informed-Neural-Networks (PINNs) have attracted considerable attention for solving mathematical models based on nonlinear PDEs; see, for example, [34, 40] and the references therein. This section addresses the numerical resolution of problem (5), and its linearized counterpart, by using Physics-Informed-Neural-Networks (PINNs).
@@ 4.1 Description of the numerical algorithm
For the sake of clarity in the exposition, let us consider the linearized system (14), with $\overline{\psi}=0,\overline{h}=1$ (which corresponds to Experiment 1 below), the nonlinear case being completely analogous. More precisely, we consider the system
$\left\{\begin{aligned} &z_{t}-z_{xx}-z=\mathds{1}_{\omega}v&&\text{in}\;\Omega_{T}:=\Omega\times(0,T),\\ &z=0&&\text{on}\;\Gamma_{T}:=\partial\Omega\times(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;\Omega,\end{aligned}\right$.
@@
[…]
@@
$\begin{array}[]{ll}\mathcal{L}\left(\boldsymbol{\theta};\mathcal{T}\right)&=w_{1}\mathcal{L}_{\text{int},PDE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int, }\Omega_{T}}\right)+w_{2}\mathcal{L}_{\text{int },ODE}\left(\boldsymbol{\theta};\mathcal{T}_{\text{int, }T}\right)\\ &+w_{3}\mathcal{L}_{\Gamma}\left(\boldsymbol{\theta};\mathcal{T}_{\Gamma}\right)\\ &+w_{4}\mathcal{L}_{t=0}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=0}\right)+w_{5}\mathcal{L}_{t=0}^{k}\left(\boldsymbol{\theta}\right)\\ &+w_{6}\mathcal{L}_{t=T}^{z}\left(\boldsymbol{\theta};\mathcal{T}_{t=T}\right)+w_{7}\mathcal{L}_{t=T}^{k}\left(\boldsymbol{\theta}\right)\\ &+w_{8}\mathcal{L}_{\hat{z}\geq 0}(\boldsymbol{\theta}),\end{array}$ (38)
@@
$\hat{z}(x,t)$$\hat{k}(t)$$\hat{v}(x,t)$$\hat{z}_{t}-\hat{z}_{xx}-\hat{z}-\mathds{1}_{\omega}\hat{v}=0$$\hat{k}^{\prime}(t)+2\mu\hat{z}_{x}(1,t)=0$$\hat{z}(-1,t)=\hat{z}(1,t)=0$$\hat{z}(x,0)=\hat{z}_{0}(x)$$\hat{k}(0)=\hat{k}_{0}$$\hat{z}(x,T)=\hat{z}_{T}(x)$$\hat{k}(T)=\hat{k}_{T}$$\min_{\theta}\mathcal{L}(\hat{z},\hat{k},\hat{v};\theta)$$\theta^{\ast}$$\hat{z}(x,t;\theta^{\ast})$$\hat{k}(t,\theta^{\ast})$$\hat{v}(x,t;\theta^{\ast})$
@@
Step 4: Training process. The final step in the PINN algorithm amounts to minimizing (38), i.e.,